# A Human Behavioral Baseline for Collective Governance in Software Projects

### v1

@@ 1 Introduction
The use of artificial intelligence systems in software project management has become increasingly salient (Hashimzai and Mohammadi, 2024). In addition to assisting individual developers, they are coordinating core management functions, including drafting pull requests, triaging issues, proposing code reviews, and enforcing release gates. As these capabilities are embedded in team tooling, decision rights migrate from human maintainers toward sociotechnical pipelines. In these pipelines, algorithms and people jointly govern workflows (Xiao et al., 2024; Wessel et al., 2022). This shift raises questions about how authority is redistributed when algorithms mediate both individual contributions and collective coordination (Crawford et al., 2023). Insights from recent work on algorithmic collective action indicate that when multiple groups interact with the same algorithmic system, their strategies can interfere in unexpected ways. A campaign that achieves near perfect success in isolation may see its efficacy drop sharply when a second group acts at the same time (Karan et al., 2025). We use algorithmic collective action to denote coordinated behavior among participants when interactions are shaped by algorithmic systems within sociotechnical platforms. In software development, the growing reality of human and algorithmic co-production leads us to ask: how might AI systems embedded in team support tools reshape governance structures, stakeholder participation, and power relations on platforms, and what options exist for steering them toward the common good (Varanasi and Goyal, 2023)?
@@
The debate about how AI systems reshape governance has outpaced empirical evidence (Delgado et al., 2023). Scholars of AI governance and participatory design warn that algorithmic infrastructures can undermine stakeholder agency, reinforce hierarchies, or reallocate decision rights (Birhane et al., 2022). The recency of AI technologies make it currently challenging to gain substantial inferences on trends of human-AI interactions in software project management. In addition, most studies examine these emergent dynamics through case studies, audits, or simulations without a historical baseline against which to evaluate the change (Margetts, 2022). As a result, claims about AI-induced shifts in governance remain speculative. Establishing a historical baseline for governance change is therefore the central objective of this paper. Before assessing whether AI systems redistribute power, narrow opportunities for participation, or enable more inclusive governance, it is necessary to understand how authority has evolved in primarily human-governed settings (Sharma et al., 2020). Our study addresses this need by constructing a large-scale longitudinal baseline of institutional change in open source project governance before the widespread adoption of AI-managed tools, providing a reference point for evaluating how future AI-mediated platforms may reshape participation and oversight. To our knowledge, this is the first large-scale longitudinal baseline of open-source governance prior to the widespread adoption of AI-managed tools.
@@
Open source software communities have been extensively studied as exemplar instances of collective knowledge work Benkler (2006); Heckman et al. (2007); Lee and Cole (2003); O’Mahony and Ferraro (2007); Schweik and English (2012); Chakraborti et al. (2024a). Importantly, they are a transparent testbed for studying governance (O’Mahony, 2005). Open source projects externalize governance in version-controlled files such as GOVERNANCE.md, which makes rules explicit, textual, and historically archived. Because governance edits are version-controlled and public, rule changes are observable at fine temporal resolution and comparable across time and projects. This is unique to OSS, which, unlike conventional organizational settings, supports systematic observation of how authority changes over time.
@@ 2 Methods
Starting with a seeded collection and filename patterns, we analyzed 710 repositories with at least one governance file at the repository root. The corpus spans 2013–2022, with governance commits recorded through June 2022 (earliest: 2013-05-09; latest: 2022-05-19). File coverage is dominated by GOVERNANCE.md, which appears in 673 out of 710 projects (94.8%), alongside 37 filename variants. The latest governance file is present for all projects, and Markdown structure is detectable in 498 repositories (70.1%), with a median of 5 sections (range: 2 to 25) Yan et al. (2023).
@@
Normalization and alignment. We preprocess governance documents by removing badges and images, converting tables to lists, normalizing headings, and stripping markup. Text is segmented into short paragraph blocks and sentences using a splitter tuned for Markdown lists. To reduce pronoun ambiguity, we apply coreference resolution while maintaining a reversible mapping to original offsets (Lee et al., 2018) (Jurafsky and Martin, ). Where headings are detectable, we record section counts across snapshots to capture the degree of governance structuring.
@@
Institutional parsing. The governance structure of each GitHub project in our corpus was extracted from its GOVERNANCE.md constitution using the Institutional Grammar (IG) framework (Ostrom, 2009) (Crawford & Ostrom, 1995; Ostrom, 2006). This framework maps the syntactic elements of policy texts to institutional primitives, first decomposing paragraphs and multi-phrase sentences into simple "institutional statements". Under the institutional grammar, an institution is treated as a bag of institutional statements. Recent NLP methods have made their automated extraction from policy text feasible Rice et al. (2021); Chakraborti et al. (2024b). Governance documents were parsed into institutional statements consisting of four linked components. An institutional statement has a Role (known in IG as ’Attribute’) when its grammatical subject is a kind of agent. Roles account for the types of actor or position recognized by the institution (e.g., "project lead," "contributor"). An Action (the ‘Aim’ in IG; syntactically the verb) identifies activities recognized by the institution as requiring governance (e.g., "commit," "assign," "review"). A Deontic captures the prescriptive force of the institutional statement, expressed through modal verbs such as "may," "should," or "must," which indicate whether an action is permitted, recommended, or required. Deontics can also be enabling ("can") or restricting ("cannot"). An Object represents the grammatical object of the rule, whether another actor or a resource that is subject to the action enacted by the statement’s role. For example, in the sentence "The technical committee must ratify the development roadmap", the Role is "technical committee," the Action is "ratify," the Object is "roadmap," and the Deontic is "must," which renders the statement obligatory.
@@
We extracted these components with the NLP4Gov toolkit (Chakraborti et al., 2024b) (Chakraborti et al., 2024a), which combines dependency parsing with semantic role labeling to parse each unitary institutional statement into its IG components. The parser emits tuples with and anchors spans and positions, enabling traceability back to the original text. Modal verbs such as may, can, should, must, and will were canonicalized into a closed set of deontic types, while role names such as maintainer, committer, reviewer, and release manager were normalized into a controlled vocabulary manually. We further manually categorized Actions into a version of the Typology of Rules adapted from the institutional analysis literature Ostrom (2009); Weible et al. (2012); Weible and Heikkila (2017); Weible et al. (2018). To test the reliability of these qualitative steps of the analysis, two authors coded the same sample of 50 Actions, which demanded the most manual categorization among the four types of institutional features. Over this sample they demonstrated a percent agreement of $82\%$ and a Cohen’s $\kappa=0.92$, over 9 labels, including a null label, strong evidence for the intersubjective validity of the chosen typology.
@@
Embedding, clustering, and metrics. Each canonical tuple is rendered as a short governance statement. We encode each governance statement with a Sentence-BERT encoder (Reimers and Gurevych, 2019) and apply BERTopic (Grootendorst, 2022) to derive semantic clusters per repository at two snapshots (initial, latest). BERTopic operates in the embedding space to form topic–like groups and uses class–based TF–IDF to label them. To ensure even clustering across all the projects’ corpus, we use the library’s standard hyperparameters without custom tuning. For structure, we report (i) richness $K$ as the number of distinct clusters per repository and (ii) normalized Shannon entropy $H$ (bits, base 2) over cluster proportions, with longitudinal change $\Delta H=H_{\text{latest}}-H_{\text{initial}}$. For drift, we compute Jensen–Shannon divergence (JSD, bits) between the aligned initial and latest cluster distributions for each repository. All repository–level estimates are aggregated with equal–weight bootstrap confidence intervals by resampling repositories with replacement.
@@ Appendix A Methods
Figure bellow presents an end-to-end pipeline that transforms raw governance documents into structured institutional insights. It begins with data normalization and pairing rules that align governance snapshots across versions. Coreference resolution reduces pronoun ambiguity, enabling more accurate attribution of roles. Semantic Role Labeling maps sentences to predicate-argument structures, identifying the underlying grammar of governance actions. These structured tuples are then embedded and clustered using BERTopic to capture governance topologies. The resulting clusters are evaluated using metrics such as entropy, Jensen–Shannon divergence, and per-project cluster counts to quantify structural diversity, prescriptiveness, and change. This modular pipeline supports scalable, interpretable analysis of institutional evolution in OSS projects.
Figure 1: End-to-end pipeline from raw governance files to structured institutional statements and analysis.

### v2

@@ 1 Introduction
The use of artificial intelligence systems in software project management has become increasingly salient [hashimzai2024integration]. In addition to assisting individual developers, they are coordinating core management functions, including drafting pull requests, triaging issues, proposing code reviews, and enforcing release gates. As these capabilities are embedded in team tooling, decision rights migrate from human maintainers toward sociotechnical pipelines. In these pipelines, algorithms and people jointly govern workflows [xiao2024generative, wessel2022quality]. This shift raises questions about how authority is redistributed when algorithms mediate both individual contributions and collective coordination [crawford2023ai]. Insights from recent work on algorithmic collective action indicate that when multiple groups interact with the same algorithmic system, their strategies can interfere in unexpected ways. A campaign that achieves near perfect success in isolation may see its efficacy drop sharply when a second group acts at the same time [karan2025algorithmic]. We use algorithmic collective action to denote coordinated behavior among participants when interactions are shaped by algorithmic systems within sociotechnical platforms. In software development, the growing reality of human and algorithmic co-production leads us to ask: how might AI systems embedded in team support tools reshape governance structures, stakeholder participation, and power relations on platforms, and what options exist for steering them toward the common good [varanasi2023currently]?
@@
The debate about how AI systems reshape governance has outpaced empirical evidence [delgado2023participatory]. Scholars of AI governance and participatory design warn that algorithmic infrastructures can undermine stakeholder agency, reinforce hierarchies, or reallocate decision rights [birhane2022power]. The recency of AI technologies make it currently challenging to gain substantial inferences on trends of human-AI interactions in software project management. In addition, most studies examine these emergent dynamics through case studies, audits, or simulations without a historical baseline against which to evaluate the change [margetts2022rethinking]. As a result, claims about AI-induced shifts in governance remain speculative. Establishing a historical baseline for governance change is therefore the central objective of this paper. Before assessing whether AI systems redistribute power, narrow opportunities for participation, or enable more inclusive governance, it is necessary to understand how authority has evolved in primarily human-governed settings [sharma2020artificial]. Our study addresses this need by constructing a large-scale longitudinal baseline of institutional change in open source project governance before the widespread adoption of AI-managed tools, providing a reference point for evaluating how future AI-mediated platforms may reshape participation and oversight. To our knowledge, this is the first large-scale longitudinal baseline of open-source governance prior to the widespread adoption of AI-managed tools.
@@
Open source software communities have been extensively studied as exemplar instances of collective knowledge work [benkler2006wealth, heckman2007emergent, 10.1287/orsc.14.6.633.24866, 10.5465/amj.2007.27169153, schweik2012internet, chakraborti2024we]. Importantly, they are a transparent testbed for studying governance [o2005nonprofit]. Open source projects externalize governance in version-controlled files such as GOVERNANCE.md, which makes rules explicit, textual, and historically archived. Because governance edits are version-controlled and public, rule changes are observable at fine temporal resolution and comparable across time and projects. This is unique to OSS, which, unlike conventional organizational settings, supports systematic observation of how authority changes over time.
@@ 2 Methods
Starting with a seeded collection and filename patterns, we analyzed 710 repositories with at least one governance file at the repository root. The corpus spans 2013–2022, with governance commits recorded through June 2022 (earliest: 2013-05-09; latest: 2022-05-19). File coverage is dominated by GOVERNANCE.md, which appears in 673 out of 710 projects (94.8%), alongside 37 filename variants. The latest governance file is present for all projects, and Markdown structure is detectable in 498 repositories (70.1%), with a median of 5 sections (range: 2 to 25) [10174145].
@@
Normalization and alignment. We preprocess governance documents by removing badges and images, converting tables to lists, normalizing headings, and stripping markup. Text is segmented into short paragraph blocks and sentences using a splitter tuned for Markdown lists. To reduce pronoun ambiguity, we apply coreference resolution while maintaining a reversible mapping to original offsets [lee-etal-2018-higher] [jurafskyspeech]. Where headings are detectable, we record section counts across snapshots to capture the degree of governance structuring.
@@
Institutional parsing. The governance structure of each GitHub project in our corpus was extracted from its GOVERNANCE.md constitution using the Institutional Grammar (IG) framework [ostrom2009understanding] (Crawford & Ostrom, 1995; Ostrom, 2006). This framework maps the syntactic elements of policy texts to institutional primitives, first decomposing paragraphs and multi-phrase sentences into simple "institutional statements". Under the institutional grammar, an institution is treated as a bag of institutional statements. Recent NLP methods have made their automated extraction from policy text feasible [10.1111/padm.12711, chakraborti2024nlp4gov]. Governance documents were parsed into institutional statements consisting of four linked components. An institutional statement has a Role (known in IG as ’Attribute’) when its grammatical subject is a kind of agent. Roles account for the types of actor or position recognized by the institution (e.g., "project lead," "contributor"). An Action (the ‘Aim’ in IG; syntactically the verb) identifies activities recognized by the institution as requiring governance (e.g., "commit," "assign," "review"). A Deontic captures the prescriptive force of the institutional statement, expressed through modal verbs such as "may," "should," or "must," which indicate whether an action is permitted, recommended, or required. Deontics can also be enabling ("can") or restricting ("cannot"). An Object represents the grammatical object of the rule, whether another actor or a resource that is subject to the action enacted by the statement’s role. For example, in the sentence "The technical committee must ratify the development roadmap", the Role is "technical committee," the Action is "ratify," the Object is "roadmap," and the Deontic is "must," which renders the statement obligatory.
@@
We extracted these components with the NLP4Gov toolkit [chakraborti2024nlp4gov] [chakraborti2024we], which combines dependency parsing with semantic role labeling to parse each unitary institutional statement into its IG components. The parser emits tuples with and anchors spans and positions, enabling traceability back to the original text. Modal verbs such as may, can, should, must, and will were canonicalized into a closed set of deontic types, while role names such as maintainer, committer, reviewer, and release manager were normalized into a controlled vocabulary manually. We further manually categorized Actions into a version of the Typology of Rules adapted from the institutional analysis literature [ostrom2009understanding, weible2012understanding, weible2017policy, weible2018understanding]. To test the reliability of these qualitative steps of the analysis, two authors coded the same sample of 50 Actions, which demanded the most manual categorization among the four types of institutional features. Over this sample they demonstrated a percent agreement of $82\%$ and a Cohen’s $\kappa=0.92$, over 9 labels, including a null label, strong evidence for the intersubjective validity of the chosen typology.
@@
Embedding, clustering, and metrics. Each canonical tuple is rendered as a short governance statement. We encode each governance statement with a Sentence-BERT encoder [reimers2019sentence] and apply BERTopic [grootendorst2022bertopic] to derive semantic clusters per repository at two snapshots (initial, latest). BERTopic operates in the embedding space to form topic–like groups and uses class–based TF–IDF to label them. To ensure even clustering across all the projects’ corpus, we use the library’s standard hyperparameters without custom tuning. For structure, we report (i) richness $K$ as the number of distinct clusters per repository and (ii) normalized Shannon entropy $H$ (bits, base 2) over cluster proportions, with longitudinal change $\Delta H=H_{\text{latest}}-H_{\text{initial}}$. For drift, we compute Jensen–Shannon divergence (JSD, bits) between the aligned initial and latest cluster distributions for each repository. All repository–level estimates are aggregated with equal–weight bootstrap confidence intervals by resampling repositories with replacement.
@@ 5 Conclusion
Acknowledgments
This research was supported by the National Science Foundation under the Growing Convergence Research Scheme (Grant No. 2020751). Additional support was provided by NSF RCN Award No. 1917908, “Coordinating and Advancing Analytical Approaches for Policy Design,” and NSF Smart and Connected Communities Award No. 2421385, “Designing Smart Environments to Augment Collective Learning and Creativity.” We thank our collaborators and colleagues for their helpful discussions and feedback.
@@ Appendix A Methods
Figure below presents an end-to-end pipeline that transforms raw governance documents into structured institutional insights. It begins with extracting initial and latest versions of GOVERNANCE.md files. Coreference resolution reduces pronoun ambiguity, enabling more accurate attribution of roles. Semantic Role Labeling maps sentences to predicate–argument structures, identifying the underlying grammar of governance actions. These structured tuples are then embedded and clustered using BERTopic to capture governance topologies. The resulting clusters are evaluated through per-project cluster counts, changes in the distribution of policy effort, and robustness checks. This modular pipeline supports scalable, interpretable analysis of institutional evolution in OSS projects.
Figure 1: Processing pipeline from raw governance files to structured institutional statements and analysis. These steps support measurement of change in count and concentration from initial to latest versions of version-controlled GOVERNANCE.md project constitutions. This diagram shows how governance text is normalized, parsed into roles, actions, and deontics, and clustered into institutional constructs.